Probabilities of Events | Charles Redmond | Observable
Author
Charles Redmond
@credmond
Comments (4)
Robert Majczyk
Oct 2, 2020
number assigned a 0 or 1 0 will be never 1 will be always
Reply
Robert Majczyk
Oct 2, 2020
3 chances that you can get a head side. .25+.25+.25=.75
Reply
Robert Majczyk
Oct 2, 2020
1/6 + 1/6 = 1/6 = .50
Reply
Robert Majczyk
Oct 2, 2020
.7 + .2 =.9
Reply
Add comment
Subscribe to notifications
© 2026 Observable, Inc.
Terms of service
Privacy policy
Product
About
Pricing
Documentation
Release notes
Contact us
Open source
Notebook Kit
Plot
D3
Community
Explore
Forum
Bluesky
YouTube
X
9
Follow
Follow
Probabilities of Events | Charles Redmond | Observable